Please help me solve these problems.
Solve 6x-0.37=5(x-0.28)
Solve the syslem by substitution.
2x+2y=0
y=4x+0
Solve by addition 5x-4y=-5
x=2y=-7
Please help me solve these problems
Thank you
MM

Answer by funmath(2933)

Hi MM,
I can help you with the first two, but the last one has a type-o. (I think the 0 in E2 of the second problem should be something else as well.) Let me know what you meant and I'll edit this.
:
Solve these problems.
:
6x-0.37=5(x-0.28)
6x-0.37=5x-1.4
6x-5x-0.37=5x-5x-1.4
x-0.37=-1.4
x-0.37+0.37=-1.4+0.37
x=-1.03
:
Solve the system by substitution.
E1: 2x+2y=0
E2: y=4x+0 ---> y=4x
Substitute 4x in for y in E1 and solve for x.
2x+2(4x)=0
2x+8x=0
10x=0
10x/10=0/10
x=0
Substitute 0 in for x in E2 and solve for y.
y=4(0)+0
y=0+0
y=0
The solution is (x,y)=(0,0)
Happy Calculating!!!
Solve by addition 5x-4y=-5
x=2y=-7